# Paper2Code: Automating Code Generation from Scientific Papers in Machine Learning

### v3

@@ Appendix F Examples output of the planning phase
Figure 16: Artifact from Step 1.1 (Overall Plan) in the planning stage of PaperCoder, generated during repository construction for the Transformer [42] (1).
@@
Figure 17: Artifact from Step 1.1 (Overall Plan) in the planning stage of PaperCoder, generated during repository construction for the Transformer [42] (2).
@@
Figure 18: Artifact from Step 1.2 (Architecture Design) in the planning stage of PaperCoder, generated during repository construction for the Transformer [42].
@@
Figure 19: Class diagram visualization generated from Step 1.2 (Architecture Design) in the planning stage of PaperCoder, generated during repository construction for the Transformer [42].
@@
Figure 20: Program call flow visualization generated from Step 1.2 (Architecture Design) in the planning stage of PaperCoder, generated during repository construction for the Transformer [42].
@@
Figure 21: Artifact from Step 1.3 (Logic Design) in the planning stage of PaperCoder, generated during repository construction for the Transformer [42].
@@
Figure 22: Artifact from Step 1.4 (Configuration File) in the planning stage of PaperCoder, generated during repository construction for the Transformer [42].
@@ A Simple Interpretable Transformer for Fine-Grained Image Classification and Analysis
Figure 23: Human Evaluation Format
Figure 24: Human Evaluation Format
Figure 25: Human Evaluation Format

### v4

@@ Appendix E Examples output of the planning phase
Figure 29: Artifact from Step 1.1 (Overall Plan) in the planning stage of PaperCoder, generated during repository construction for the Transformer (Vaswani et al., 2017) (1/2).
@@
Figure 30: Artifact from Step 1.1 (Overall Plan) in the planning stage of PaperCoder, generated during repository construction for the Transformer (Vaswani et al., 2017) (2/2).
@@
Figure 31: Artifact from Step 1.2 (Architecture Design) in the planning stage of PaperCoder, generated during repository construction for the Transformer (Vaswani et al., 2017).
@@
Figure 32: Class diagram visualization generated from Step 1.2 (Architecture Design) in the planning stage of PaperCoder, generated during repository construction for the Transformer (Vaswani et al., 2017).
@@
Figure 33: Program call flow visualization generated from Step 1.2 (Architecture Design) in the planning stage of PaperCoder, generated during repository construction for the Transformer (Vaswani et al., 2017).
@@
Figure 34: Artifact from Step 1.3 (Logic Design) in the planning stage of PaperCoder, generated during repository construction for the Transformer (Vaswani et al., 2017).
@@
Figure 35: Artifact from Step 1.4 (Configuration File) in the planning stage of PaperCoder, generated during repository construction for the Transformer (Vaswani et al., 2017).
@@ A Simple Interpretable Transformer for Fine-Grained Image Classification and Analysis
Figure 36: Human Evaluation Guideline (1/3)
Figure 37: Human Evaluation Guideline (2/3)
Figure 38: Human Evaluation Guideline (3/3)